# Machine-learned models for magnetic materials

### v1

@@ VI Modified Siamese model
with $N=N_{\mathrm{train}}\times N_{\mathrm{bias}}$, $M=5$, and the constant frequency nodes centered at positions $\Omega_{i}=\{10^{3},10^{4.75},10^{6.5},10^{8.25},10^{10}\}$. Now the loss function for the modified Siamese NN model is the following:
@@ VII Discussion and Conclusions
In the first approach to the problem, we defined a network that predicts the parameters directly (containing the encoder part only) and compares them with the parameters used to generate a given characteristic. Such training was strictly supervised by using the encoder loss, defined as
@@
The sum in Eq. 12 goes over the training samples and parameters: $\hat{y}_{ij}$ means the $i$-th parameter ($L_{i}$ or $R_{i}$) predicted for $j$-th sample, $Z_{j}$, whereas $y_{ij}$ represents ground truth parameters – we already know them from the generation process. However, this approach simply failed because it was not possible to effectively train a neural model that would generalize beyond a limited set of characteristics being some narrow part of the training data. Much better results came from the unsupervised training where the autoencoder model is simply given a variety of characteristics and has to attempt to reconstruct them. The model proposed in our scheme is able to generalize well, which is shown in Table I where the validation error reaches a few percent for the training set with a huge variety of characteristics. Consequently, the NN model is able to correctly fit parameters for previously unseen measured test data. At the same time, it is worth noting that the diversity of the training set fulfilled a kind of regularization, ensuring that the NN does not overfit a particular behavior.
@@
We also observe the important effect that imposing continuity of the predicted parameters as a function of the bias current using the Siamese network, i.e., forcing the transitions to be physical, causes the fit to individual characteristics in the family to improve significantly.
The NN model has also a robust tendency to parameter selection. In the experiment we have shown that too large model leads to a fit in which some components did not carry a significant contribution, thus neglecting these components effectively reduces the found analytical model while keeping its error equally small.
@@ VIII Methods
Data generation was performed with the usage of Pytorch library [35] on GPU, which significantly speedups the generation process from a few hours (when using CPU) to a dozen seconds (GPU). NN model (encoder and decoder part) was implemented using Pytorch with an automatic differentiation engine (autograd) used to track the gradient of the loss with respect to the model parameters during the backpropagation stage. The training of the single model on the GPU Nvidia A6000 took about 1 hour.
@@ IX Acknowledgments
The authors would like to acknowledge the financial support of Project No. [anonymized] granted by [anonymized].

### v2

@@ VI Modified Siamese model
with $N=N_{\mathrm{train}}\times N_{\mathrm{bias}}$, $M=5$, and the constant frequency nodes centered at positions $\Omega_{i}=\{10^{3},10^{4.75},10^{6.5},10^{8.25},10^{10}\}$ Hz. The values of $\Omega_{i}$ have been chosen to evenly cover the frequency range of the measurement data. Moreover, including an extra point above this range further improved the fitting. Now the loss function for the modified Siamese NN model is the following:
@@ VII Discussion and Conclusions
In the first approach to the problem, we defined a network that predicts the parameters directly (containing the encoder part only) and compares them with the parameters used to generate a given characteristic. Such training was strictly supervised by using the encoder loss, defined as:
@@
The sum in Eq. 12 goes over the training samples and parameters: $\hat{y}_{ij}$ means the $i$-th parameter ($L_{i}$ or $R_{i}$) predicted for $j$-th sample, $Z_{j}$, whereas $y_{ij}$ represents ground truth parameters – we already know them from the generation process. However, this approach simply failed because it was not possible to effectively train a neural model that would generalize beyond a limited set of characteristics being some narrow part of the training data. Much better results came from the unsupervised training where the AE model is simply given a variety of characteristics and has to attempt to reconstruct them. The model proposed in our scheme is able to generalize well, which is shown in Table I where the validation error reaches a few percent for the training set with a huge variety of characteristics. Consequently, the NN model is able to correctly fit parameters for previously unseen measured test data. At the same time, it is worth noting that the diversity of the training set (see Fig.2) fulfilled a kind of regularization, ensuring that the NN does not overfit any particular behavior.
In the proposed approach, the decoder is needed only to estimate the loss function ($\mathcal{L}_{\mathrm{dec}}$), that is minimized during the training. The approach where a large component of NN model (in our case, the decoder) is used only to estimate the error of the predictive part (encoder) in ML is not unusual. It is characteristic, for example, for generative adversarial networks (GANs)[49], where the discriminator is trained to estimate the error of the generator, or in neural style transfer models [50], where the pretrained NN determines so-called style or content losses.
The proposed scheme for fitting parameters of lumped model is general and in principle can be applied to any analytical model. The decoder introduces analytical formula which parameters are predicted by the encoder. The only requirement is that the decoder formula, which will pass the fitting error when the encoder is trained by backpropagation, should be differentiable.
@@
In classical methods for finding equivalent circuit elements, there is difficulty in capturing the correlation between individual $Z(f)$ characteristics for different DC-bias currents [41], and they are usually only able to match individual characteristics [51, 52]. This results in discontinuous parameters (here, the ladder elements L, R) as a function of the DC current. The Siamese network proposed here solves this problem. Vector fitting, as a standard method for finding transfer function parameters, not only has the problem of ensuring continuous parameters as a function of DC current but also cannot guarantee that the identified transfer function is a positive real function[53]. Using neural networks allows for the imposition of arbitrary constraints and boundary conditions, including enforcing the physical behavior of the fitted parameters.
The proposed method exhibits a robust tendency to parameter selection. Although the initially suggested complexity of analytical model was $M=7$, the NN network identified some components as having negligible contributions (as shown in Fig. 5(b)), effectively reducing the complexity $M\rightarrow{}4$. Consequently, in the subsequent experiment (Fig. 5(d)), NN successfully predicted a simplified analytical model with $M=5$ while maintaining a comparably low error.
One way to further improve the presented model is to eliminate the arbitrarily determined grid of $\Omega_{i}$ (Eq. 10) by adding a term in loss function that penalizes the characteristic frequencies of the individual LR ladder elements coming too close to each other. This will result in a more uniform distribution of these frequencies without imposing their positions from above.
@@ VIII Methods
Data generation was performed with the usage of Pytorch library [54] on GPU (RTX A6000 48GB GDDR6), which significantly speedups the generation process from a few hours (when using CPU - 64 cores 128 threads Ryzen Threadripper 3990X) to a dozen seconds (GPU). NN model (encoder and decoder part) was implemented using Pytorch with an automatic differentiation engine (autograd) used to track the gradient of the loss with respect to the model parameters during the backpropagation stage. The training of the single model on the GPU Nvidia RTX A6000 took about 1 hour.
@@ IX Acknowledgments
The authors would like to acknowledge the financial support of Project No. 2019/34/E/ST7/00187 granted by National Science Centre, Poland (NCN).